You have Tk. 1,00,000 to invest. If you can earn 9% interest, about how long does it take for

your Tk. 1,00,000 investment to grow to Tk. 2,00,000?


1
Expert's answer
2021-06-01T03:29:02-0400

"2,00,000-1,00,000=1,00,000\times 0.09 \times t \\

1,00,000=9,000t \\

t = \frac{1,00,000}{9,000}=11.11 \; years"